OpenStudy (anonymous):

A gas company has the following rate schedule for natural gas usage in single-family residences: Monthly service charge $8.80 Per therm service charge 1st 25 therms $0.6686/therm Over 25 therms $0.85870/therm

OpenStudy (anonymous):

What is the charge for using 25 therms in one month? Show your work.

OpenStudy (anonymous):

@q12157

OpenStudy (anonymous):

What is the charge for using 45 therms in one month? Show your work.

OpenStudy (anonymous):

well the first 25 therms is $0.6686/therm

OpenStudy (anonymous):

So the first 25 therms costs $(0.6686*25) in total.

OpenStudy (anonymous):

since you are buying 45 therms, there are 20 (as 45-25=20) therms that arent covered by the initial price

OpenStudy (anonymous):

so the next 20 therms cost $0.85870/therm

OpenStudy (anonymous):

The total cost of these therms is $(20*0.8587)

OpenStudy (anonymous):

SO the total cost of all 45 therms is...

OpenStudy (anonymous):

$(0.6686*25)+$(20*0.8587)

OpenStudy (anonymous):

which is equal to?

OpenStudy (anonymous):

I think you can calculate the rest :p if you need help just ask

OpenStudy (anonymous):

25 therms cost $16.715 45 therms cost $33.889 you should check my answers x3

OpenStudy (anonymous):

i got this for the 25 therms 8.8 + 25*0.6686 = 16.715+8.8 = 25.52

OpenStudy (anonymous):

Oh yeah you need to add the monthly service charge... good job

OpenStudy (anonymous):

just add 8.8 to my answers :P

OpenStudy (anonymous):

25 therms cost $16.715+8.8=? 45 therms cost $33.889+8.8=?